# PAS-SE: Personalized Auxiliary-Sensor Speech Enhancement for Voice Pickup in Hearables

### v1

@@ 1 Introduction
Hearable devices with one or more microphones can be used to capture the user’s own voice (i.e., target speech), but due to environmental noise and interfering talkers, the captured speech signal needs to be enhanced to facilitate communication. Although single-channel speech enhancement (SE) approaches based on deep neural networks can reduce environmental noise [1, 2, 3, 4], they tend to struggle with removing interfering talkers from the mixture while preserving the user’s voice. Modern hearables often include microphones placed inside or near the user’s occluded ear canal, typically employed in active noise reduction systems [5]. These in-ear microphones are acoustically shielded from environmental noise and interfering talkers by the device. At the same time, the user’s voice is also picked up by the in-ear microphone, predominantly through body-conduction [6, 7]. These two effects lead to a substantial benefit in terms of signal-to-noise ratio (SNR) of the user’s voice at the in-ear microphone. However, this signal cannot be directly used for voice communication due to time- and user-varying band-limitation, nonlinearities and distortions introduced by body conduction [6, 7], and undesired additive body-produced noises [8]. Other auxiliary body-conduction sensors, such as accelerometers, have similar properties and trade-offs as in-ear microphones [9].
@@
Experimental evaluations carried out on two openly available in-ear microphone datasets demonstrate that interfering speech can be considerably reduced with either PSE or AS-SE, while a combination of both, which we term personalized auxiliary-sensor speech enhancement (PAS-SE), leads to further improvements. In particular, PAS-SE personalized with in-ear enrollments led to the best results, generalizing both within and across datasets tested.
@@ 3 System architecture
which takes the magnitude of the noisy STFT coefficients of one or more microphones as input. The architecture consists of a unidirectional long short-term memory (LSTM) layer operating across the frequency dimension of the input (F-LSTM) with 512 hidden units, a causal, unidirectional LSTM layer operating across the time dimension (T-LSTM) with 128 hidden units, a linear layer, and a $\tanh$ activation function. The output is a magnitude mask $M(k,l)$, which is applied to the noisy outer microphone signal $Y_{o}(k,l)$, i.e. $\hat{S}_{o}(k,l)=M(k,l)\cdot Y_{o}(k,l)$. The SE system (with one input) has 1.384 M parameters, and the AS-SE system has 1.386 M parameters.
We add a conditioning mechanism to the architecture to investigate personalization for our PSE and PAS-SE systems by introducing a speaker encoder branch based on the time-domain Speakerbeam architecture [15] using the authors’ implementation in the Asteroid toolkit [22]. An utterance $\tilde{Y}_{\text{enroll}}(k,l)$ from the same user but excluding the input utterance (either recorded at the outer or in-ear microphone) is used as input to the speaker encoder branch. The branch consists of a learnable filterbank encoder, which is followed by a 1D-convolutional block and temporal averaging to obtain a 128-dimensional speaker embedding vector $\mathbf{e}$. A dense layer is then used to match the dimension of the output of the F-LSTM layer of the main enhancement network. The output of the F-LSTM layer is multiplied with the output of the dense layer from the speaker encoder branch. The speaker encoder consists of 1.810 M parameters.
@@ 4.1 Datasets
The same procedure for mixing own voice, noise, and interferers is carried out for both datasets. For each utterance during training, there is a 75% probability of environmental noise being added. If added, own voice and environmental noise are mixed to an SNR randomly selected from an uniform distribution between [-10, 10] dB (defined at the outer microphone). The corresponding scaling factor is applied to the in-ear microphone signals as well, to realistically preserve SNR differences between the noisy outer and acoustically shielded in-ear microphone. The same procedure is applied to interferers, with an independent 75% probability. This ensures the system is trained on a combination of user’s clean speech, noisy speech, noisy speech with interferers, and speech with only interferers. Since Vibravox does not contain an isolated interfering speaker partition, we randomly select from the set of speakers in the training partition disjoint from the target speaker as the interferer at each step.
@@ 4.2 In-ear noise and interferer training configurations
To facilitate training AS-SE and PAS-SE systems without recorded interferer signals (as in Vibravox), we explore various configurations of incorporating in-ear signal noise and interferer components during training as shown in Table 1. The components in (1) are either included using recorded signals (✓) or not (✗) during training, or the in-ear microphone component is approximated by the corresponding outer microphone component attenuated using a random factor $a\in[0.001,1]$ from a uniform distribution (empirically chosen).
@@ 4.3 Experimental setup
The experiments are conducted at a sampling rate of 16 kHz, using an STFT framework with a frame length of 32 ms and a frame shift of 16 ms, where a square-root Hann window is used both in analysis and synthesis. Training is carried out with a batch size of 8 and an example length of 3 seconds (randomly selected clip). Mean-variance-normalization based on clean speech training subset statistics is applied to each input channel independently. For personalized systems, a single disjoint utterance $\tilde{S}$ from the same talker was randomly selected as the clean enrollment speech $\tilde{Y}_{\text{enroll}}=\tilde{S}_{\text{enroll}{\{o,i\}}}$ (OM or IM enrollment) for each training example.
@@ 5 Results
For evaluation, we use the scale-invariant signal-to-distortion ratio (SI-SDR) [28], wideband PESQ [29], and ESTOI [30] between the target and estimated clean speech signal at the outer microphone [^1] Audio examples & paper resources: https://bose.github.io/passe/. . Results are averaged across the entire test set. For evaluation, target speech is either mixed with noise (N), interfering voices (V) or both (N+V). To evaluate robustness against noisy enrollment speech in Section 5.3, clean enrollment utterances are mixed with noise $\tilde{N}$, so $\tilde{Y}_{\text{enroll}}=\tilde{S}_{\text{enroll}{\{o,i\}}}+\tilde{N}_{\{o,i\}}$, and the performance of the system is evaluated only adding interfering speech (V) to the target speech.
@@ 5.1 In-domain evaluation
Table 3 shows the results for noise reduction by AS-SE and PAS-SE systems trained with different configurations of noise and interferer in-ear components evaluated on the Vibravox dataset. When no in-ear noise or interferer components are used (A), performance is poor, but including in-ear noise in training (B, C, D) yields large improvements, highlighting the importance of modeling the noise leakage to the in-ear microphone in the training. As expected for noise reduction, different configurations of interferer in-ear components do not lead to substantial performance differences between configurations (B), (C), and (D). Similarly, added personalization (PAS-SE) does not yield improvements.
@@ 5.2 Cross-dataset generalization
Overall, the performance trends indicate that the AS-SE systems require the use of some level of interferer signals during training (C, D) to generalize to an unseen auxiliary-sensor array. Cross-domain performance improves more when combined with enrollment-based personalization (PAS-SE), nearing or even outperforming the corresponding matched system trained on in-domain data (OL).
@@ 5.3 Personalization with noisy enrollment signals
Fig. 2: Cross-dataset interferer reduction performance (V) achieved by SE, PSE, AS-SE and PAS-SE systems at different enrollment utterance SNRs (-$\infty$: only noise, no speech, $\infty$: clean speech).

### v2

@@ 1 Introduction
Hearable devices with one or more microphones can be used to capture the user’s own voice (i.e., target speech), but due to environmental noise and interfering talkers, the captured speech signal needs to be enhanced to facilitate communication. Although single-channel speech enhancement (SE) approaches based on deep neural networks can reduce environmental noise [1, 2, 3, 4], they tend to struggle with removing interfering talkers from the mixture while preserving the user’s voice. Modern hearables often include microphones placed inside or near the user’s occluded ear canal, typically employed in active noise reduction systems [5]. These in-ear microphones are acoustically shielded from environmental noise and interfering talkers by the device. At the same time, the user’s voice is also picked up by the in-ear microphone, predominantly through body conduction [6, 7]. These two effects lead to a substantial benefit in terms of signal-to-noise ratio (SNR) of the user’s voice at the in-ear microphone. However, this signal cannot be directly used for voice communication due to time- and user-varying band-limitation, nonlinearities and distortions introduced by body conduction [6, 7], and undesired additive body-produced noises [8]. Other auxiliary body-conduction sensors, such as accelerometers, have similar properties and trade-offs as in-ear microphones [9].
@@
Experimental evaluations carried out on two openly available in-ear microphone datasets demonstrate that interfering speech can be considerably reduced with either PSE or AS-SE, while a combination of both, which we term personalized auxiliary-sensor speech enhancement (PAS-SE), leads to further improvements. In particular, PAS-SE personalized with in-ear enrollments leads to the best results, generalizing both within and across datasets tested. Code and audio examples are available online [^1] https://bose.github.io/passe/. .
@@ 3 System architecture
which takes the magnitude of the noisy STFT coefficients of one or more microphones as input. The architecture consists of a unidirectional long short-term memory (LSTM) layer operating across the frequency dimension of the input (F-LSTM) with 512 hidden units, a causal unidirectional LSTM layer operating across the time dimension (T-LSTM) with 128 hidden units, a linear layer, and a $\tanh$ activation function. The output is a magnitude mask $M(k,l)$, which is applied to the noisy outer microphone signal $Y_{o}(k,l)$, i.e. $\hat{S}_{o}(k,l)=M(k,l)\cdot Y_{o}(k,l)$. The SE system (with one input) has 1.384 M parameters, and the AS-SE system has 1.386 M parameters.
We add a conditioning mechanism to the architecture to investigate personalization for our PSE and PAS-SE systems by introducing a speaker encoder branch based on the time-domain SpeakerBeam architecture [15] using the authors’ implementation in the Asteroid toolkit [22]. An utterance $\tilde{Y}_{\text{enroll}}(k,l)$ from the same user, excluding the input utterance (either recorded at the outer or in-ear microphone), is used as input to the speaker encoder branch. The branch consists of a learnable filterbank encoder, which is followed by a 1-D convolutional block and temporal averaging to obtain a 128-dimensional speaker embedding vector $\mathbf{e}$. A dense layer is then used to match the dimension of the output of the F-LSTM layer of the main enhancement network. The output of the F-LSTM layer is multiplied with the output of the dense layer from the speaker encoder branch. The speaker encoder consists of 1.810 M parameters.
@@ 4.1 Datasets
The same procedure for mixing own voice, noise, and interferers is carried out for both datasets. For each utterance during training, there is a 75% probability of environmental noise being added. If added, own voice and environmental noise are mixed to an SNR randomly selected from a uniform distribution between [-10, 10] dB (defined at the outer microphone). The corresponding scaling factor is applied to the in-ear microphone signals as well, to realistically preserve SNR differences between the noisy outer and acoustically shielded in-ear microphone. The same procedure is applied to interferers, with an independent 75% probability. This ensures the system is trained on a combination of user’s clean speech, noisy speech, noisy speech with interferers, and speech with only interferers. Since Vibravox does not contain an isolated interfering speaker partition, we randomly select from the set of speakers in the training partition disjoint from the target speaker as the interferer at each step.
@@ 4.2 In-ear noise and interferer training configurations
To facilitate training of AS-SE and PAS-SE systems without recorded interferer signals (as in Vibravox), we explore various configurations of incorporating in-ear signal noise and interferer components during training as shown in Table 1. The components in (1) are either included using recorded signals (✓) or not (✗) during training, or the in-ear microphone component is approximated by the corresponding outer microphone component attenuated using a random factor $a\in[0.001,1]$ from a uniform distribution.
@@ 4.3 Experimental setup
The experiments are conducted at a sampling rate of 16 kHz, using an STFT framework with a frame length of 32 ms and a frame shift of 16 ms, where a square-root Hann window is used both in analysis and synthesis. Training is carried out with a batch size of 8 and an example length of 3 seconds (randomly selected clip). Mean-variance normalization based on clean speech training subset statistics is applied to each input channel independently. For personalized systems, a single disjoint utterance $\tilde{S}$ from the same talker was randomly selected as the clean enrollment speech $\tilde{Y}_{\text{enroll}}=\tilde{S}_{\text{enroll}{\{o,i\}}}$ (OM or IM enrollment) for each training example.
@@ 5 Results
For evaluation, we use the scale-invariant signal-to-distortion ratio (SI-SDR) [28], wideband PESQ [29], and ESTOI [30] between the target and estimated clean speech signal at the outer microphone. Results are averaged across the entire test set. For evaluation, target speech is either mixed with noise (N), interfering voices (V) or both (N+V). To evaluate robustness against noisy enrollment speech in Section 5.3, clean enrollment utterances are mixed with noise $\tilde{N}$, so $\tilde{Y}_{\text{enroll}}=\tilde{S}_{\text{enroll}{\{o,i\}}}+\tilde{N}_{\{o,i\}}$, and the performance of the system is evaluated only adding interfering speech (V) to the target speech.
@@ 5.1 In-domain evaluation
Table 3 shows the results for noise reduction by AS-SE and PAS-SE systems trained with different configurations of noise and interferer in-ear components evaluated on the Vibravox dataset. When no in-ear noise or interferer components are used (A), performance is poor, but including in-ear noise in training (B, C, D) yields large improvements, highlighting the importance of modeling the noise leakage to the in-ear microphone during training. As expected for noise reduction, different configurations of interferer in-ear components do not lead to substantial performance differences between configurations (B), (C), and (D). Similarly, added personalization (PAS-SE) does not yield improvements.
@@ 5.2 Cross-dataset generalization
These performance trends indicate that the AS-SE systems require the use of some level of interferer signals during training (C, D) to generalize to an unseen auxiliary-sensor array. Cross-domain performance improves more when combined with enrollment-based personalization (PAS-SE), nearing or even outperforming the corresponding matched system trained on in-domain data (OL).
@@ 5.3 Personalization with noisy enrollment signals
Fig. 2: Cross-dataset interferer reduction performance (V) achieved by SE, PSE, AS-SE, and PAS-SE systems at different enrollment utterance SNRs (-$\infty$: only noise, no speech, $\infty$: clean speech).